OpenStudy (anonymous):

Eight tickets to a play cost $96. How much do 3 tickets cost? A. $12 B. $36 C. $48 D. $60

OpenStudy (anonymous):

B each ticket cost 12$

OpenStudy (secret-ninja):

If it takes 96 dolar for 8, then you just divide 96 by 8. Once you do, you get 12. It is 12 dollars per ticket, but you need to know the value of 3 tickets, so you do 12 x 3 = 36 dollars for 3 tickets.

OpenStudy (secret-ninja):

So your answer would be B. $36

OpenStudy (secret-ninja):

I was kind of confused for a moment there... lol

OpenStudy (texaschic101):

proportion time 8/96 = 3/x --- 8 tickets to 96 = 3 ticket to x cross multiply 8x = 96(3) 8x = 288 x = 288/8 x = 36 3 tickets cost $ 36

OpenStudy (secret-ninja):

@lordoftanks If this helped, please click the "Best Response" button. :)